OpenStudy (anonymous):

[9.01] and [9.02] For y = x2 − 4x + 3, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation. Show all work and use complete sentences to receive full credit.

OpenStudy (scorcher219396):

If there's no negative in front of the x^2 (provided y is not negative too) then it opens upwards If the parabola opens upwards, then the vertex will be a minimum To find the vertex, put your equation in graphing form by completing the square (tell me if that step confuses you) You'll get y=(x-2)^2-1 which is y=a(x-h)^2+k, and the vertex is at (h,k) So this one is at (2,-1) To find the x intercepts, set the y=0 and then factor the x's, you'll get 0=(x-3)(x-1) and set each of those =0, you'll get x=1,3 and those are your x intercepts Describe the graph... Opens upwards, is normal width, drops below the x axis 2 and moves left 1

OpenStudy (anonymous):

thank you so much, can you help me with another one?

OpenStudy (anonymous):

[9.06] Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 3, – 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x − ___ Numerical Answers Expected!

OpenStudy (scorcher219396):

Since it gives you the x intercepts, you can think back to how you would have found the x intercepts You would have gotten two things factored that when you set them equal to 0 you got -8 and 2, so what would those two factored things be? (x+8) and (x-2) Now if you work back even more, and multiply those two back together to unfactor, you'll get the original equation, which would be x^2+6x-16